Hello out there,I need some help with some problems,the first one is,Solve:3y/2=7+y/4.The next one is,Find the product:8a/a+7 times 2a+14/16a.The third one is,Find the quotient:b^2/b-6 divided b/b-6.And the last one is,Find the quotient:5cd divided 10c/d,thanks to whoever can help.
Answer by stanbon(75887)

the first one is,Solve:
3y/2=7+y/4.
Multiply thru by 4 to get:
6y = 28 + y
5y = 28
y = 5 3/8
----------------------
The next one is,Find the product:
8a/a+7 times 2a+14/16a.
[8a/(a+7)]
Cancel where you can:
Cancel: 8a, and (a+7)
=[1/1]
= 1
==========
Cheers,
Stan H.






The third one is,Find the quotient:b^2/b-6 divided b/b-6.
And the last one is,Find the quotient:5cd divided 10c/d
-----------------------